OpenStudy (anonymous):

Which one is a rule for the nth term of the sequence 3, 15, 75, 375...?

OpenStudy (anonymous):

I believe domain times 5?

OpenStudy (anonymous):

nth term would be 5 times (n-1)th term

OpenStudy (anonymous):

im confused...haha

OpenStudy (anonymous):

so this is a recursive sequence where every next term depends on previous term

OpenStudy (anonymous):

the 2 possible answers a_n=3(5)^n-1 and a_n=5(3)^n-1

OpenStudy (anonymous):

if its n-1. there are also answers with n+1

OpenStudy (anonymous):

umm you would have to check that then